OpenStudy (anonymous):

Juliana has created the function f(x) = 3x + 2/ 4 to represent the cost of texting on her current plan, where x represents the number of texts. Juliana discovers that using the inverse function to solve for x = 24, she can predict how many texts she can use for $24. Explain to Juliana how to accomplish this, using complete sentences.

OpenStudy (anonymous):

So far, I replaced the f(x) with y and then I switched y and x \[y = \frac{ 3x+2 }{ 4 }\]

OpenStudy (anonymous):

then I switched the y and x and then multiplied both sides by 4 \[(4)x = \frac{ 3y +2 }{ 4 } (4)\]

OpenStudy (anonymous):

then, after canceling out the fraction, I subtracted 2 on both sides --> 4x - 2 = 3y. Now i'm not sure if I should replace the x value with 24 or if i'm doing this all wrong?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Nurali

hero (hero):

Divide both sides by 3 now. Then change \(y\) to \(f^{-1}(x)\)

OpenStudy (anonymous):

Where would I divide 3? so far I got 4x - 2 = 3y

hero (hero):

\[\frac{4x - 2}{3} = y\]

OpenStudy (anonymous):

Wait, since it started out as f(x) = 3x +2 over 4, to make it an inverse you change the equation? how do you know the equation is like that though?

hero (hero):

How do you know it is the inverse? Well, there is one way to know for sure. Compose both functions. If the result you get is x, then you have the correct inverse.

OpenStudy (anonymous):

I still don't understand, what do I do now that I have 4x - 2/ 3 =y? do I multiply by 3 on both sides?

hero (hero):

You have isolated y, which is the goal after swapping x and y. Now you replace y with \(f^{-1}(x)\)

hero (hero):

\(f^{-1}(x)\) stands for the inverse of \(f(x)\)

OpenStudy (anonymous):

So you are telling me that the answer is 4x - 2/ 3 = y?? is there no more?

OpenStudy (anonymous):

what about x = 24

hero (hero):

I didn't say you were done with the word problem. I said, you are done finding the inverse once you replace y with \(f^{-1}(x)\)

hero (hero):

You can continue solving the word problem now that you have the inverse.

OpenStudy (anonymous):

So would I now fill in x with 24?

OpenStudy (anonymous):

If I filled in x with 24 my answer would be 31.3333 and I don't think that 31.3333 is correct right?

hero (hero):

It is correct, but of course you round to 31

OpenStudy (anonymous):

Really? so it would be 31m= F-1(x)? Would I say that her answer is 31?

OpenStudy (anonymous):

*31= f-1(x)

hero (hero):

You would say, the number of text she can make for $24 is 31 texts.

OpenStudy (anonymous):

Ahh ok, thank you very very much :)